# Estimation of Optimal Causal Bounds via Covariate-Assisted Optimal Transport

### v1

@@ Appendix C Proofs of Section 3
Here, the first equation is due to Proposition 2.2.
For (Term A): we highlight that we consider the optimal coupling $\pi\in\mathcal{P}(\mathcal{Y}^{2}\times\mathcal{Z}^{2})$ that is attained in the adapted Wasserstein distance $\mathds{W}_{\textup{bc}}(\mu_{Y(1),Z},\nu_{Y(1),Z})$ to connect $\mu_{Z}$ and $\nu_{Z}$, then
[…]
@@
$\leq$ $L_{h}\int\mathds{W}_{1}\left(\mu_{Y(0)}^{z},\mu_{Y(0)}^{z^{\prime}}\right)+\mathds{W}_{1}\left(\mu_{Y(1)}^{z},\nu_{Y(1)}^{z^{\prime}}\right){\rm d}\pi_{Z,Z^{\prime}}(z,z^{\prime})$
$\leq$ $L_{h}\int L_{Z}\|z-z^{\prime}\|_{2}+\mathds{W}_{1}\left(\mu_{Y(1)}^{z},\nu_{Y(1)}^{z^{\prime}}\right){\rm d}\pi_{Z,Z^{\prime}}(z,z^{\prime})$
$\leq$ $L_{h}(L_{Z}\vee 1)\mathds{W}_{\textup{bc}}(\mu_{Y(1),Z},\nu_{Y(1),Z}).$
Here, the second inequality is due to Lemma C.2; the third inequality is due to Assp. 3.4; the last inequality is due to the definition of $\pi$, which achieves the optimality in $\mathds{W}_{\textup{bc}}(\mu_{Y(1),Z},\nu_{Y(1),Z})$.
@@
$\leq$ $L_{h}\int\mathds{W}_{1}(\mu_{Y(0)}^{z},\nu_{Y(0)}^{z}){\rm d}\nu_{Z}(z),$
where the last inequality is due to Lemma C.2.
Now, we consider the optimal coupling $\tilde{\pi}\in\mathcal{P}(\mathcal{Y}^{2}\times\mathcal{Z}^{2})$ that is attained in $\mathds{W}_{\textup{bc}}(\mu_{Y(0),Z},\nu_{Y(0),Z})$, then
$\begin{split}&\int\mathds{W}_{1}(\mu_{Y(0)}^{z},\nu_{Y(0)}^{z}){\rm d}\nu_{Z}(z)\\ =&\int\mathds{W}_{1}(\mu_{Y(0)}^{z^{\prime}},\nu_{Y(0)}^{z^{\prime}}){\rm d}\tilde{\pi}(z,z^{\prime})\\ =&\int\mathds{W}_{1}(\mu_{Y(0)}^{z^{\prime}},\nu_{Y(0)}^{z^{\prime}})-\mathds{W}_{1}(\mu_{Y(0)}^{z^{\prime}},\mu_{Y(0)}^{z}){\rm d}\tilde{\pi}(z,z^{\prime})+\int\mathds{W}_{1}(\mu_{Y(0)}^{z^{\prime}},\mu_{Y(0)}^{z}){\rm d}\tilde{\pi}(z,z^{\prime})\\ \leq&\int\mathds{W}_{1}(\nu_{Y(0)}^{z^{\prime}},\mu_{Y(0)}^{z}){\rm d}\tilde{\pi}(z,z^{\prime})+L_{Z}\int\left\|{z-z^{\prime}}\right\|_{{2}}{\rm d}\tilde{\pi}(z,z^{\prime})\\ \leq&(L_{Z}\vee 1)\mathds{W}_{\textup{bc}}(\mu_{Y(0),Z},\nu_{Y(0),Z}).\end{split}$ (12)
Here, the last to second inequality is due to Lemma C.2 and Assp. 3.4; the last inequality is due to the definition of $\tilde{\pi}$.
Therefore, we get
(Term $(\text{Term B})\leq L_{h}(L_{Z}\vee 1)\mathds{W}_{\textup{bc}}(\mu_{Y(0),Z},\nu_{Y(0),Z}).$
@@
$\leq$ $L_{h}(L_{Z}\vee 1)(\mathds{W}_{\textup{bc}}(\mu_{Y(0),Z},\nu_{Y(0),Z})+\mathds{W}_{\textup{bc}}(\mu_{Y(1),Z},\nu_{Y(1),Z})).$
Replace $\nu$ by the measures from the sequence of $(\nu_{n},n\geq 1)$, we get the desired result. ∎
Appendix D Proofs of Section 4
D.1 Proof of Theorem 4.2
The proof of Theorem 4.2 requires several useful results from [backhoff2022estimating]. Since our definition of adapted empirical distribution is slightly different from that of [backhoff2022estimating], to make the paper self-contained, we include proofs of the lemmas in Section E.
Remark D.1 (Notation).
We introduce the following notation, which is aligned with [backhoff2022estimating]. For notational convenience, in this proof and the proofs of the associated lemmas and corollaries, we write $\hat{\bm{\mu}}^{z_{0}}_{Y(0)}$ to represent $\hat{\bm{\mu}}^{z_{0}}_{Y(0),n}$ and $\hat{\bm{\mu}}^{z_{1}}_{Y(1)}$ to represent $\hat{\bm{\mu}}^{z_{1}}_{Y(1),m}$. We also write $\hat{\pi}$ to represent $\hat{\pi}_{n,m}$. The constants, if not specified explicitly, are independent of $n,m$ but may depend on $d_{Z},d_{Y},M$ (Assp. 3.2), $L_{Z},L_{h}$.
Let $\Phi_{r}^{n}$ be the small cubes that partition $[0,1]^{d_{Z}}$ as defined in Definition 4.1, i.e.,
@@ Appendix D Proofs of Section 4
Lemma D.1 (Variance of $\mathds{W}_{h}$).
Under Assp. 3.3, the following holds.
@@
Lemma D.2 (Gap between $\hat{\bm{\mu}}$ and $\hat{\mu}$).
Under Assp. 3.1, we have
@@
Lemma D.3 (Gap between $\hat{\bm{\mu}}$ and $\mu$).
Under Assp. 3.1 and 3.4, we have
@@
Lemma D.4 ([backhoff2022estimating, Lemma 3.4]).
@@
    Under Assp. 3.1, and 4.1, $\mathbb{E}[\mathds{W}_{1}(\hat{\mu}_{Z(0)},\mu_{Z})]\leq C_{1}R_{d_{Z}}(n)$.
@@
Here, $R_{d}(\cdot)$ is defined in (10), $C_{1}$ is a constant that depends on $d_{Z}$, and $C_{2}$ is a constant that depends on $d_{Y},M$ (Assp. 3.2).
Proof of Theorem 4.2.
@@
For the integrand, we apply Lemma C.2 and get
$\left|\mathds{W}_{h}(\hat{\bm{\mu}}^{z_{0}}_{Y(0)},\hat{\bm{\mu}}^{z_{1}}_{Y(1)})-\mathds{W}_{h}(\mu^{z_{0}}_{Y(0)},\mu^{z_{1}}_{Y(1)})\right|\leq L_{h}(\mathds{W}_{1}(\hat{\bm{\mu}}^{z_{0}}_{Y(0)},\mu^{z_{0}}_{Y(0)})+\mathds{W}_{1}(\hat{\bm{\mu}}^{z_{1}}_{Y(1)},\mu^{z_{1}}_{Y(1)})).$ (13)
@@
For (Term B), again, we apply Lemma C.2 to the integrand and get
$\left|\mathds{W}_{h}(\mu^{z_{0}}_{Y(0)},\mu^{z_{1}}_{Y(1)})-\mathds{W}_{h}(\mu^{z_{0}}_{Y(0)},\mu^{z_{0}}_{Y(1)})\right|\leq$ $L_{h}\mathds{W}_{1}(\mu^{z_{0}}_{Y(1)},\mu^{z_{1}}_{Y(1)})\leq L_{h}L_{Z}\left\|{z_{0}-z_{1}}\right\|_{{2}},$
where the second inequality is due to Assp. 3.4. Therefore, we have
@@
By Lemma C.2 and Assp. 3.4, the mapping $z\mapsto\mathds{W}_{h}(\mu^{z}_{Y(0)},\mu^{z}_{Y(1)})$ is $(2L_{h}L_{Z})$-Lipschitz continuous. Then, by Lemma D.2,
@@
By Assp. 4.1, Lemma D.1 and Markov’s inequality,
@@
For (Term C), we apply Assp. 4.2. Finally, combine (Term C) and (Term D), and we arrive at
[…]
@@
Then, apply Lemma D.4, we get
@@
$R_{d_{Y}}(\cdot)$ is defined in (10), and $C_{2}$ is a constant that depends on $d_{Y},M$ (Assp. 3.2).
@@
Remark D.2 (Choice of $\hat{\pi}_{n,m}$ in (8)).
When $\hat{\pi}_{n,m}$ is the optimal coupling of the marginals, by Lemma D.2 and the triangle inequality (Lemma A.1),
@@
Therefore, $\hat{\pi}_{n,m}$ in (8) satisfies Assp. 4.2 when $r=\frac{1}{d_{Z}+2\vee d_{Y}}$.
Remark D.3 (Proof of Remark 4.1).
If $\mu_{Z(1)}\neq\mu_{Z}$, then for $r=\frac{1}{d_{Z}+2\vee d_{Y}}$, according to the derivation in Remark D.2, we have
@@
$\leq$ $2\sqrt{d_{Z}}(n^{-r}+m^{-r})+\mathbb{E}\left[\mathds{W}_{1}(\hat{\mu}_{Z(0)},\mu_{Z})\right]+\mathbb{E}\left[\mathds{W}_{1}(\hat{\mu}_{Z(1)},\mu_{Z(1)})\right]+\mathds{W}_{1}(\mu_{Z(1)},\mu_{Z})$
$\leq$ $C^{\prime}(n\wedge m)^{-r}+\epsilon.$
Plug this condition back to the derivation in the proof of Theorem 4.2, specifically (15), we get the desired result.
D.2 Proof of Theorem 4.1
Lemma D.5 (Convergence of adapted empirical distribution).
Under Assp. 3.1 - 3.4, then almost surely,
$\lim_{n\rightarrow\infty}\mathds{W}_{\textup{bc}}(\hat{\bm{\mu}}_{Y(0),Z(0),n},\mu_{Y(0),Z})=\lim_{m\rightarrow\infty}\mathds{W}_{\textup{bc}}(\hat{\bm{\mu}}_{Y(1),Z(1),m},\mu_{Y(1),Z})=0.$
Proof of Theorem 4.1.
The proof strategy of Theorem 4.1 is similar to that of Theorem 4.2. By (14), we have
$|\hat{V}_{m,n}-V_{\textup{c}}|$
@@
By Lemma D.5, we get (Term A) converges to zero almost surely.
For (Term B), by Assp. 3.13.4, $(z_{0},z_{1})\mapsto\mathds{W}_{h}(\mu^{z_{0}}_{Y(0)},\mu^{z_{1}}_{Y(1)})$ is a bounded continuous function. Indeed, by Assp. 3.4 and Lemma C.2,
$\mathds{W}_{h}(\mu^{z_{0}}_{Y(0)},\mu^{z_{1}}_{Y(1)})-\mathds{W}_{h}(\mu^{z_{0}^{\prime}}_{Y(0)},\mu^{z_{1}^{\prime}}_{Y(1)})\leq$ $L_{h}(\mathds{W}_{1}(\mu^{z_{0}}_{Y(0)},\mu^{z_{0}^{\prime}}_{Y(0)})+\mathds{W}_{1}(\mu^{z_{1}}_{Y(1)},\mu^{z_{1}^{\prime}}_{Y(1)}))$
@@
Further, by Assp. 3.1, we get that $\mathds{W}_{h}(\mu^{z_{0}}_{Y(0)},\mu^{z_{1}}_{Y(1)})$ is bounded on $\mathcal{Z}$.
@@
D.3 Proof of Theorem 4.3
Lemma D.6 (TV bound).

### v2

@@ E.2 Proof of Theorem 2
Here, the first equation is due to Proposition 2.
By Assp. 4-5, $h$ is continuous and supported on a compact domain, then $h$ is uniformly continuous. Therefore, for any $\epsilon>0$, there is a constant $\widetilde{C}(\epsilon)$ such that for any $y,y^{\prime}\in\mathcal{Y}$,
$|h(y)-h(y^{\prime})|\leq\widetilde{C}(\epsilon)\left\|{y-y^{\prime}}\right\|_{{2}}+\epsilon.$
Thus, Lemma 5 is applicable and for any $(\mu,\nu,\mu^{\prime},\nu^{\prime})\in\mathcal{P}(\mathcal{Y})$,
$\left|\mathds{W}_{h}(\mu,\nu)-\mathds{W}_{h}(\mu^{\prime},\nu^{\prime})\right|\leq\widetilde{C}(\epsilon)\left(\mathds{W}_{1}(\mu,\mu^{\prime})+\mathds{W}_{1}(\nu,\nu^{\prime})\right)+2\epsilon.$ (11)
Similarly, by Lemma 4, we have
$\mathds{W}_{1}\left(\mu_{Y(k)}^{z_{0}},\mu_{Y(k)}^{z_{1}}\right)\leq C(\epsilon)\|z_{0}-z_{1}\|_{2}+\epsilon.$ (12)
For (Term A): we highlight that we consider the optimal coupling $\pi\in\mathcal{P}(\mathcal{Y}^{2}\times\mathcal{Z}^{2})$ that is attained in the adapted Wasserstein distance $\mathds{W}_{\textup{a}}(\mu_{Y(1),Z},\nu_{Y(1),Z})$ to connect $\mu_{Z}$ and $\nu_{Z}$, then for $\epsilon,\epsilon^{\prime}>0$,
$(\text{Term A})=$ $\left|\int\mathds{W}_{h}\left(\mu_{Y(0)}^{z},\mu_{Y(1)}^{z}\right){\rm d}\mu_{Z}(z)-\int\mathds{W}_{h}\left(\mu_{Y(0)}^{z},\nu_{Y(1)}^{z}\right){\rm d}\nu_{Z}(z)\right|$
@@
$\leq$ $\widetilde{C}(\epsilon)\int\mathds{W}_{1}\left(\mu_{Y(0)}^{z},\mu_{Y(0)}^{z^{\prime}}\right)+\mathds{W}_{1}\left(\mu_{Y(1)}^{z},\nu_{Y(1)}^{z^{\prime}}\right){\rm d}\pi_{Z,Z^{\prime}}(z,z^{\prime})+2\epsilon$
$\leq$ $\widetilde{C}(\epsilon)\int(C(\epsilon^{\prime})\|z-z^{\prime}\|_{2}+\epsilon^{\prime})+\mathds{W}_{1}\left(\mu_{Y(1)}^{z},\nu_{Y(1)}^{z^{\prime}}\right){\rm d}\pi_{Z,Z^{\prime}}(z,z^{\prime})+2\epsilon$
$\leq$ $\widetilde{C}(\epsilon)(C(\epsilon^{\prime})\vee 1)\mathds{W}_{\textup{a}}(\mu_{Y(1),Z},\nu_{Y(1),Z})+2\epsilon+\widetilde{C}(\epsilon)\epsilon^{\prime}.$
Here, the second inequality is due to (11); the third inequality is due to (12); the last inequality is due to the definition of $\pi$, which achieves the optimality in $\mathds{W}_{\textup{a}}(\mu_{Y(1),Z},\nu_{Y(1),Z})$.
@@
$(\text{Term B})=$ $\left|\int\mathds{W}_{h}\left(\mu_{Y(0)}^{z},\nu_{Y(1)}^{z}\right){\rm d}\nu_{Z}(z)-\int\mathds{W}_{h}\left(\nu_{Y(0)}^{z},\nu_{Y(1)}^{z}\right){\rm d}\nu_{Z}(z)\right|$
$\leq$ $\widetilde{C}(\epsilon)\int\mathds{W}_{1}(\mu_{Y(0)}^{z},\nu_{Y(0)}^{z}){\rm d}\nu_{Z}(z)+2\epsilon,$
where the last inequality is due to (11).
Now, we consider the optimal coupling $\tilde{\pi}\in\mathcal{P}(\mathcal{Y}^{2}\times\mathcal{Z}^{2})$ that is attained in $\mathds{W}_{\textup{a}}(\mu_{Y(0),Z},\nu_{Y(0),Z})$, then
$\begin{split}&\int\mathds{W}_{1}(\mu_{Y(0)}^{z},\nu_{Y(0)}^{z}){\rm d}\nu_{Z}(z)\\ =&\int\mathds{W}_{1}(\mu_{Y(0)}^{z^{\prime}},\nu_{Y(0)}^{z^{\prime}}){\rm d}\tilde{\pi}(z,z^{\prime})\\ =&\int\mathds{W}_{1}(\mu_{Y(0)}^{z^{\prime}},\nu_{Y(0)}^{z^{\prime}})-\mathds{W}_{1}(\mu_{Y(0)}^{z^{\prime}},\mu_{Y(0)}^{z}){\rm d}\tilde{\pi}(z,z^{\prime})+\int\mathds{W}_{1}(\mu_{Y(0)}^{z^{\prime}},\mu_{Y(0)}^{z}){\rm d}\tilde{\pi}(z,z^{\prime})\\ \leq&\int\mathds{W}_{1}(\nu_{Y(0)}^{z^{\prime}},\mu_{Y(0)}^{z}){\rm d}\tilde{\pi}(z,z^{\prime})+\int\mathds{W}_{1}(\mu_{Y(0)}^{z^{\prime}},\mu_{Y(0)}^{z}){\rm d}\tilde{\pi}(z,z^{\prime})\\ \leq&\int\mathds{W}_{1}(\nu_{Y(0)}^{z^{\prime}},\mu_{Y(0)}^{z}){\rm d}\tilde{\pi}(z,z^{\prime})+C(\epsilon^{\prime})\int\left\|{z-z^{\prime}}\right\|_{{2}}{\rm d}\tilde{\pi}(z,z^{\prime})+\epsilon^{\prime}\\ \leq&(C(\epsilon^{\prime})\vee 1)\mathds{W}_{\textup{a}}(\mu_{Y(0),Z},\nu_{Y(0),Z})+\epsilon^{\prime}.\end{split}$ (13)
Here, the first inequality is due to the triangle inequality of the metric $\mathds{W}_{1}$; the second inequality is due to (12); the last inequality is due to the definition of $\tilde{\pi}$.
Therefore, we get
@@
$\leq$ $\widetilde{C}(\epsilon)(C(\epsilon^{\prime})\vee 1)(\mathds{W}_{\textup{a}}(\mu_{Y(0),Z},\nu_{Y(0),Z})+\mathds{W}_{\textup{a}}(\mu_{Y(1),Z},\nu_{Y(1),Z}))+2(2\epsilon+\widetilde{C}(\epsilon)\epsilon^{\prime}).$
Replace $\nu$ by the measures from the sequence of $(\nu_{n},n\geq 1)$ and send $\epsilon^{\prime},\epsilon\rightarrow 0$, we get the desired result. ∎
Appendix F Proofs of Section 4
Notation. In this section and the following, to simplify the notation, we will let $n_{0}=n,n_{1}=m$, and $\xi_{0}=w,\xi_{1}=\xi$.
F.1 Proof of Theorem 4
The proof of Theorem 4 requires several useful results from [3]. Since our definition of adapted empirical distribution is slightly different from that of [3], to make the paper self-contained, we include proofs of the lemmas in Section H.
Remark 3 (Notation).
We introduce the following notation, which is aligned with [3]. For notational convenience, in this proof and the proofs of the associated lemmas and corollaries, we write $\hat{\bm{\mu}}^{z_{0}}_{Y(0)}$ to represent $\hat{\bm{\mu}}^{z_{0}}_{Y(0),n}$ and $\hat{\bm{\mu}}^{z_{1}}_{Y(1)}$ to represent $\hat{\bm{\mu}}^{z_{1}}_{Y(1),m}$. We also write $\hat{\pi}$ to represent $\hat{\pi}_{n,m}$. The constants, if not specified explicitly, are independent of $n,m$ but may depend on $d_{Z},d_{Y}$, $L_{Z},L_{h}$.
Let $\Phi_{r}^{n}$ be the small cubes that partition $0,1]^{d_{Z}}$ as defined in Definition 2, i.e.,
@@ F.1 Proof of Theorem 4
Lemma 6 (Variance of $\mathds{W}_{h}$).
Under Assp. 7, the following holds.
@@
Lemma 7 (Gap between $\hat{\bm{\mu}}$ and $\hat{\mu}$).
Under Assp. 4, we have
@@
Lemma 8 (Gap between $\hat{\bm{\mu}}$ and $\mu$).
Under Assp. 4 and 8, we have
@@
Lemma 9 ([3, Lemma 3.4]).
@@
    Under Assp. 4 and 6, $\mathbb{E}[\mathds{W}_{1}(\hat{\mu}_{Z(0)},\mu_{Z})]\leq C_{1}R_{d_{Z}}(n)$.
@@
Here, $R_{d}(\cdot)$ is defined in (8), $C_{1}$ is a constant that depends on $d_{Z}$, and $C_{2}$ is a constant that depends on $d_{Y}$.
Proof of Theorem 4.
@@
For the integrand, we apply Lemma 5 and get
$).\displaystyle\left|\mathds{W}_{h}(\hat{\bm{\mu}}^{z_{0}}_{Y(0)},\hat{\bm{\mu}}^{z_{1}}_{Y(1)})-\mathds{W}_{h}(\mu^{z_{0}}_{Y(0)},\mu^{z_{1}}_{Y(1)})\right|\leq L_{h}(\mathds{W}_{1}(\hat{\bm{\mu}}^{z_{0}}_{Y(0)},\mu^{z_{0}}_{Y(0)})+\mathds{W}_{1}(\hat{\bm{\mu}}^{z_{1}}_{Y(1)},\mu^{z_{1}}_{Y(1)}$)). (14)
@@
For (Term B), again, we apply Lemma 5 to the integrand and get
@@
where the second inequality is due to Assp. 8. Therefore, we have
$(\textup{Term B})$
@@
[…]
@@
By Assp. 6, Lemma 6 and Markov’s inequality,
@@
For (Term C), we apply Assp. 9. Finally, combine (Term C) and (Term D), and we arrive at
$\mathbb{E}[\textup{(Term B)}]\leq$ $C^{\prime}(m\wedge n)^{-r}.$
@@
Then, apply Lemma 9, we get
@@
$R_{d_{Y}}(\cdot)$ is defined in (8), and $C_{2}$ is a constant that depends on $d_{Y}$ and $\max_{y\in\mathcal{Y}}\left\|{y}\right\|_{{2}}$.
@@
Remark 4 (Choice of $\hat{\pi}_{n,m}$ in (6)).
When $\hat{\pi}_{n,m}$ is the optimal coupling of the marginals, by Lemma 7 and the triangle inequality (Lemma 1),
@@
Therefore, $\hat{\pi}_{n,m}$ in (6) satisfies Assp. 9 when $r=\frac{1}{d_{Z}+2\vee d_{Y}}$.
F.2 Proof of Theorem 5
Proof of Theorem 5.
If the generating distribution of $Z(0),Z(1)$ are not equal to $\mu_{Z}$, i.e., $\mu_{Z(0)}\neq\mu_{Z}$ and $\mu_{Z(1)}\neq\mu_{Z}$, then for $r=\frac{1}{d_{Z}+2\vee d_{Y}}$, according to the derivation in Remark 4, we have
@@ F.2 Proof of Theorem 5
$\leq$ $2\sqrt{d_{Z}}(n^{-r}+m^{-r})+\mathbb{E}\left[\mathds{W}_{1}(\hat{\mu}_{Z(0)},\mu_{Z(0)})\right]+\mathbb{E}\left[\mathds{W}_{1}(\hat{\mu}_{Z(1)},\mu_{Z(1)})\right]+\mathds{W}_{1}(\mu_{Z(0)},\mu_{Z})+\mathds{W}_{1}(\mu_{Z(1)},\mu_{Z})$
$\leq$ $C^{\prime}(n\wedge m)^{-r}+2\epsilon.$
Plug this condition back to the derivation in the proof of Theorem 4, specifically (16), we get the desired result. ∎
F.3 Proof of Theorem 3
Lemma 10 (Convergence of adapted empirical distribution).
Under Assp. 4 - 8, then almost surely,
$\lim_{n\rightarrow\infty}\mathds{W}_{\textup{a}}(\hat{\bm{\mu}}_{Y(0),Z(0),n},\mu_{Y(0),Z})=\lim_{m\rightarrow\infty}\mathds{W}_{\textup{a}}(\hat{\bm{\mu}}_{Y(1),Z(1),m},\mu_{Y(1),Z})=0.$
Proof of Theorem 3.
The proof strategy of Theorem 3 is similar to that of Theorem 4. By (15), we have
$|\hat{V}_{n,m}-V_{\textup{c}}|$
@@ F.3 Proof of Theorem 3
$\text{(Term A)}\leq(L_{Z}\vee 1)(\mathds{W}_{\textup{a}}(\hat{\bm{\mu}}_{Y(0),Z(0),n},\mu_{Y(0),Z})+\mathds{W}_{\textup{a}}(\hat{\bm{\mu}}_{Y(1),Z(1),m},\mu_{Y(1),Z})).$
By Lemma 10, we get (Term A) converges to zero almost surely.
@@
[…]
@@
[…]
@@
F.4 Proof of Theorem 6
Lemma 11 (TV bound).